# Lecture Notes on Spectral Independence and Bases of a Matroid: Local-to-Global and Trickle-Down from a Markov Chain Perspective

### v3

@@ 5.4 Key Technical Lemma
The following lemma will be the key tool in the inductive proof of Theorem 4. It relates the Dirichlet form for the up-down walk $\mathrm{P}_{k}^{\wedge}$ at level $k$ with the down-up walk $\mathrm{P}_{k}^{\vee}$ at level $k$. Recall that the spectral gap of the $\mathrm{P}_{k}^{\wedge}$ is equal to the spectral gap of $\mathrm{P}_{k+1}^{\wedge}$ (see Lemma 5.4). Hence, the lemma is in essence an inductive statement as it relates the chain $\mathrm{P}_{k}^{\wedge}$ to $\mathrm{P}_{k}^{\vee}$, rather than $\mathrm{P}_{k+1}^{\vee}$. In this manner the lemma will immediately yield an inductive proof of the Random Walk Theorem.
@@
At an intuitive level, the reader can understand the factor $k/(k+1)$ because in the LHS the down operator is from level $k+1$ to $k$ and hence a transition occurs with probability $1/(k+1)$, whereas in the RHS the down operator is from level $k$ to $k-1$ and hence occurs with probability $1/k$.
We will utilize the following two claims to prove Lemma 5.5. The first claim captures the local-to-global nature of the process. In particular, for the chain $\mathrm{P}_{k}^{\wedge}$ we can express its Dirichlet form in terms of local walks for pinnings at level $k-1$. Recall that the Dirichlet form for a chain $P$ is the local variation of a functional $f$ over pairs of states connected by a transition of $P$.
@@
$=\frac{k}{k+1}\gamma_{k-1}\mathcal{E}_{\mathrm{P}_{k}^{\vee}}(f).$ by Claim 5.7
@@
$\pi_{k}(\eta\cup a)=\frac{1}{{n\choose k}}\mu(\eta\cup a)=\frac{k}{(n-k+1)}\pi_{k-1}(\eta)\mu_{\eta}(a)=k\pi_{k-1}(\eta)\pi_{\eta,1}(a).$ (20)
@@
$=\frac{k}{k+1}\sum_{\eta\in\mathcal{P}_{k-1}}\pi_{k-1}(\eta)\mathcal{E}_{Q_{\eta}}(f_{\eta}).$
@@
$\sum_{\eta\in\mathcal{P}_{k-1}}\pi_{k-1}(\eta)\mathrm{Var}_{\pi_{\eta,1}}(f_{\eta})$ $=\frac{1}{2}\sum_{\eta\in\mathcal{P}_{k-1}}\pi_{k-1}(\eta)\sum_{\begin{subarray}{c}a,b\in V\setminus S\times\{0,1\}:\\ \eta:S\rightarrow\{0,1\}\end{subarray}}\pi_{\eta,1}(\eta\cup a)\pi_{\eta,1}(\eta\cup b)(f_{\eta}(a)-f_{\eta}(b))^{2}$
@@
$=\mathcal{E}_{\mathrm{P}_{k}^{\vee}}(f).$
5.5 Inductive Proof of Random Walk Theorem: Proof of Theorem 4
We can now prove Eq. 18, which implies the Random Walk Theorem (Theorem 4).
Proof of Theorem 4.
@@ 5.5 Inductive Proof of Random Walk Theorem: Proof of Theorem 4
6 Optimal Spectral Gap for Linear-Blocks
The goal of this section is to prove that when the system is $\eta$-spectrally independent, then the relaxation time of the Glauber dynamics is $O(n)$, thereby proving Theorem 2 of [CLV21]. Our proof follows the same approach as in [CLV21].
@@ 6.1 Uniform Block Dynamics
Consider the following dynamics, parameterized by $10<\alpha<$ 1, which we will refer to as the $\alpha n$-uniform block dynamics. The dynamics updates $\alpha n$ random vertices $S$ in each step. Assume $\alpha n$ is an integer. From $Y_{t}\in\Omega$, the transitions $Y_{t}\rightarrow Y_{t+1}$ of the $\alpha n$-uniform block dynamics are as follows:
@@
    Select $\alpha n$ random vertices, chosen uniformly at random from all (${n\choose\alpha n}$ vertices. Let $S$ denote the chosen set.
@@ 6.2 Improved Random Walk Theorem
The Random Walk Theorem is not strong enough to establish fast mixing of the $\alpha n$-uniform block dynamics, we will need the following improved local-to-global theorem which was presented by [CLV21, Theorem 5.4] and independently [GM20, Theorem 4]. For simplicity we only present the spectral gap version, the stronger version for entropy contraction is presented in [CLV21, GM20].
@@
Theorem 5 (Improved RW Theorem [CLV21, Theorem A.9],[GM20, Theorem 4]).
@@
Before proving the above theorem let us first compare it to the Random Walk Theorem (Theorem 4) stated earlier. Looking at Eq. 24 for $k=n$ (or (25) for $\ell=n-1$) we have:
@@
In contrast, Theorem 4 has $12\gamma_{j}-$ 1 replaced by $\gamma_{j}$ in the lower bound on the spectral gap of $\mathrm{P}_{n}^{\vee}$.
@@ 6.3 Fast Mixing of Uniform Block Dynamics
Applying Theorem 5 we establish fast mixing of the $\alpha n$-uniform block dynamics.
@@
For all $\eta>0$ and $10<\alpha<$ 1 there exists a constant $C=C(\eta,\alpha)$ such that if the system is $\eta$-spectral independence then the spectral gap of the $\alpha n$-uniform block dynamics is $\gamma(\mathrm{P}_{n,(1-\alpha)n}^{\vee})\geq C$.
@@
$\Gamma_{(1-\alpha/2)n}=\prod_{i=0}^{(1-\alpha/2)n-2}(2\gamma_{i}-1)\geq\prod_{i=0}^{(1-\alpha/2)n-1}\left(1-\frac{2\eta}{n-i-1}\right)\geq\left(1-\frac{4\eta}{\alpha n}\right)^{n}\geq\exp(-8\eta/\alpha),$ (26)
@@ 6.4 Shattering
The first step is to look at the properties of the updated vertices in a step of the block dynamics. In particular, [CLV21] showed that a random subset of $\alpha n$ vertices is “shattered” in the sense that the expected size of each component is $O(1)$ when $\alpha<1/(6\Delta)$ where $\Delta$ is the maximum degree of the graph $G$. The shattering occurs because $\alpha\Delta<1$ and hence this corresponds to a sub-critical branching process.
@@
[CLV21, Lemma 4.3] For a graph $G=(V,E)$ of maximum degree $\Delta$, for $\alpha>0$, choose a random subset $S\subset V$ where $|S|=\alpha n$. Then for every integer $k\geq 1$, for every $v\in V$,
$\mathrm{Prob}\left(|T_{v}|=k\right)\leq\alpha(6\Delta\alpha)^{k-1},$
where $T_{v}$ is the component containing $v$ in $\mathcal{C}_{v}$, which is the induced subgraph on $S$.
@@
See the proof of Lemma 4.3 on page 22 of [CLV21]. ∎
6.5 Optimal Relaxation Time of Glauber: Proof of Theorem 2
Using the above shattering result (Lemma 6.2) with the fast mixing result for the $\alpha n$-uniform block dynamics (Lemma 6.1), we can prove optimal upper bound on the relaxation time of the Glauber dynamics and hence prove Theorem 2.
Proof of Theorem 2.
@@ 6.5 Optimal Relaxation Time of Glauber: Proof of Theorem 2
$\mathcal{E}_{\mathrm{P_{Gl}}}(f)$ $=\frac{1}{2}\sum_{\tau\in\Omega}\sum_{\sigma\in\Omega}\mu(\tau)\mathrm{P_{Gl}}(\tau,\sigma)(f(\tau)-f(\sigma))^{2}$
$=\sum_{\tau\in\Omega}\mu(\tau)\frac{1}{n}\sum_{v\in V}\left[\frac{1}{2}\sum_{\begin{subarray}{c}\sigma\in\Omega:\\ \sigma(V\setminus\{v\})=\tau(V\setminus\{v\})\end{subarray}}\mu(\sigma)(f(\tau)-f(\sigma))^{2}\right]$ (30)
@@
$=\frac{1}{C}\frac{1}{{n\choose\alpha n}}\sum_{\begin{subarray}{c}S\subset V:\\ |S|=\alpha n\end{subarray}}\sum_{\tau\in\Omega}\mu(\tau)\mathcal{E}_{\mathrm{P_{HB}^{\tau}}(S)}(f)$ (see below)
$=\frac{1}{C}\frac{1}{{n\choose\alpha n}}\sum_{\begin{subarray}{c}S\subset V:\\ |S|=\alpha n\end{subarray}}\sum_{\tau\in\Omega}\mu(\tau)\mathrm{Var}_{S}[F\mid\tau]$ by Eq. 27
$=\frac{1}{C}\frac{1}{{n\choose\alpha n}}\sum_{\begin{subarray}{c}S\subset V:\\ |S|=\alpha n\end{subarray}}\sum_{\tau\in\Omega}\mu(\tau)\sum_{T\in\mathcal{C}_{S}}\mathrm{Var}_{T}[F\mid\tau]$ by independence of $T,T^{\prime}\in\mathcal{C}_{S}$
$\leq\frac{1}{C}\frac{1}{{n\choose\alpha n}}\sum_{\begin{subarray}{c}S\subset V:\\ |S|=\alpha n\end{subarray}}\sum_{\tau\in\Omega}\mu(\tau)\sum_{T\in\mathcal{C}_{S}}C^{\prime}|T|^{\eta+1}\sum_{v\in T}\mathcal{E}_{\mathrm{P^{\tau}_{Gl}}(T)}(f)$ by Theorem 1, (see below)
$=\frac{1}{C}\frac{1}{{n\choose\alpha n}}\sum_{\begin{subarray}{c}S\subset V:\\ |S|=\alpha n\end{subarray}}\sum_{\tau\in\Omega}\mu(\tau)\sum_{T\in\mathcal{C}_{S}}C^{\prime}|T|^{\eta+1}\frac{1}{|T|}\sum_{v\in T}\mathrm{Var}_{v}[F\mid\tau]$ by Eq. 31
$=\frac{1}{C}\sum_{\tau\in\Omega}\mu(\tau)\sum_{v\in V}\mathrm{Var}_{v}[F\mid\tau]\sum_{k=1}^{\alpha n}\mathrm{Prob}\left(|T_{v}|=k\right)\times C^{\prime}k^{\eta-1}$ rearranging
@@
$=\frac{n}{C^{\prime\prime}}\mathcal{E}_{\mathrm{P_{Gl}}}(f)$ by Eq. 31.
In the second line we are implementing the $(n,(1-\alpha)n)$ down-up dynamics by first choosing the subset $S$ of size $\alpha n$ for update and then applying the heat-bath dynamics on this set $S$ with those vertices outside $S$ fixed to $\tau$ chosen from the Gibbs distribution $\mu$. And in the fifth line where we are applying Theorem 1, note we are considering the Glauber dynamics which only operates on the vertices in this component $T$, where the vertices outside $T$ have the fixed configuration $\tau$.
This proves that the spectral gap of the Glauber dynamics is $\geq C^{\prime\prime}/n$ for a constant $C^{\prime\prime}=C^{\prime\prime}(\alpha,\Delta)$. Hence, the relaxation time of the Glauber dynamics is $O(n)$ as claimed. ∎
6.6 Improved Random Walk Theorem: Proof of Theorem 5
In this section we will prove the improved Random Walk Theorem presented in Theorem 5. The proof will involve statements comparing Dirichlet forms and variance on different levels and hence we will need to define the projection of a function $f$ on lower levels.
@@ 6.6 Improved Random Walk Theorem: Proof of Theorem 5
We defer the proofs of Lemmas 6.3 and 6.4 to Section 6.7, and we proceed with the proof of Theorem 5.
Proof of Theorem 5.
We begin by proving Eq. 24 which states that the spectral gap of $\mathrm{P}_{k}^{\vee}$ is at least $\Gamma_{k-1}/(\sum_{i=0}^{k-1}\Gamma_{i}$ where $\Gamma_{i}=\prod_{j=0}^{i-1}(2\gamma_{j}-1)$. We will then use this result for the one-level down-up walk to derive the more general statement in Theorem 5 for the down-up walk $\mathrm{P}_{n,\ell}^{\vee}$.
@@
$=\Gamma_{k}\mathcal{E}_{\mathrm{P}_{k+1}^{\vee}}(f^{(k+1)})+\Gamma_{k}\mathrm{Var}_{\pi_{k}}(f^{(k)})$ since $\mbox{since }\Gamma_{k}=\gamma_{k-1}\Gamma_{k-1}$
@@
Finally, applying Lemma 6.3 with $i=n,j=\ell$, we have that Eq. 37 is equivalent to Eq. 35, which completes the proof of Theorem 5. ∎

### v4

@@ 5.4 Key Technical Lemma
The following lemma will be the key tool in the inductive proof of Theorem 4.1. It relates the Dirichlet form for the up-down walk $\mathrm{P}_{k}^{\wedge}$ at level $k$ with the down-up walk $\mathrm{P}_{k}^{\vee}$ at level $k$. Recall that the spectral gap of the $\mathrm{P}_{k}^{\wedge}$ is equal to the spectral gap of $\mathrm{P}_{k+1}^{\wedge}$ (see Lemma 5.4). Hence, the lemma is in essence an inductive statement as it relates the chain $\mathrm{P}_{k}^{\wedge}$ to $\mathrm{P}_{k}^{\vee}$, rather than $\mathrm{P}_{k+1}^{\vee}$. In this manner the lemma will immediately yield an inductive proof of the Random Walk Theorem.
@@
At an intuitive level, the reader can understand the factor $k/(k+1)$ because in the LHS the down operator is from level $k+1$ to $k$ and hence a transition occurs with probability $1/(k+1)$, whereas in the RHS the down operator is from level $k$ to k1k-1 and hence occurs with probability $1/k$.
We will utilize the following two claims to prove Lemma 5.5. The first claim captures the local-to-global nature of the process. In particular, for the chain $\mathrm{P}_{k}^{\wedge}$ we can express its Dirichlet form in terms of local walks for pinnings at level k1k-1. Recall that the Dirichlet form for a chain $P$ is the local variation of a functional $f$ over pairs of states connected by a transition of $P$.
@@
$=\frac{k}{k+1}\gamma_{k-1}\mathcal{E}_{\mathrm{P}_{k}^{\vee}}(f),$ by Claim 5.7
as claimed. ∎
@@
$\pi_{k}(\eta\cup a)=\frac{1}{\binom{n}{k}}\mu(\eta\cup a)=\frac{k}{(n-k+1)}\pi_{k-1}(\eta)\mu_{\eta}(a)=k\pi_{k-1}(\eta)\pi_{\eta,1}(a).$ (47)
@@
$=\frac{k}{k+1}\sum_{\eta\in\mathcal{P}_{k-1}}\pi_{k-1}(\eta)\mathcal{E}_{Q_{\eta}}(f_{\eta}).\qed$
@@
$\sum_{\eta\in\mathcal{P}_{k-1}}\pi_{k-1}(\eta)\mathrm{Var}_{\pi_{\eta,1}}(f_{\eta})$
$=\frac{1}{2}\sum_{\eta\in\mathcal{P}_{k-1}}\pi_{k-1}(\eta)\sum_{\begin{subarray}{c}a,b\in V\setminus S\times\{0,1\}:\\ \eta:S\rightarrow\{0,1\}\end{subarray}}\pi_{\eta,1}(\eta\cup a)\pi_{\eta,1}(\eta\cup b)(f_{\eta}(a)-f_{\eta}(b))^{2}$
@@
$=\mathcal{E}_{\mathrm{P}_{k}^{\vee}}(f).\qed$
5.5 Inductive Proof of Random Walk Theorem: Proof of Theorem 4.1
We can now prove Eq. 45, which implies the Random Walk Theorem (Theorem 4.1).
Proof of Theorem 4.1.
@@
6 Optimal Mixing Time for Glauber from Spectral Independence
The goal of this section is to prove that when the system is $\eta$-spectrally independent, then the relaxation time of the Glauber dynamics is $O(n)$, thereby proving Theorem 1.5 of [CLV21a]. Our proof follows the same approach as in [CLV21a]. In Section 6.8 we comment on how the proof changes to obtain Theorem 1.8 establishing an optimal $O(n\log{n})$ mixing time, instead of relaxation time.
@@ 6.1 Uniform Block Dynamics
Consider the following dynamics, parameterized by $0<\alpha<1$, which we will refer to as the $\alpha n$-uniform block dynamics. The dynamics updates $\alpha n$ random vertices $S$ in each step. Assume $\alpha n$ is an integer. From $Y_{t}\in\Omega$, the transitions $Y_{t}\rightarrow Y_{t+1}$ of the $\alpha n$-uniform block dynamics are as follows:
@@
    Select $\alpha n$ random vertices, chosen uniformly at random from all $\binom{n}{\alpha n}$ vertices. Let $S$ denote the chosen set.
@@ 6.2 Improved Random Walk Theorem
The Random Walk Theorem is not strong enough to establish fast mixing of the $\alpha n$-uniform block dynamics, we will need the following improved local-to-global theorem which was presented by [CLV21a, Theorem 5.4] and independently [GM20, Theorem 4]. For simplicity we only present the spectral gap version, the stronger version for entropy contraction is presented in [CLV21a, GM20].
@@
Theorem 6.1 (Improved RW Theorem [CLV21a, Theorem A.9],[GM20, Theorem 4]).
@@
Before proving the above theorem let us first compare it to the Random Walk Theorem (Theorem 4.1) stated earlier. Looking at Eq. 51 for $k=n$ (or (52) for $\ell=n-1$) we have:
@@
In contrast, Theorem 4.1 has $2\gamma_{j}-1$ replaced by $\gamma_{j}$ in the lower bound on the spectral gap of $\mathrm{P}_{n}^{\vee}$.
@@ 6.3 Fast Mixing of Uniform Block Dynamics
Applying Theorem 6.1 we establish fast mixing of the $\alpha n$-uniform block dynamics.
@@
For all $\eta>0$ and $0<\alpha<1$ there exists a constant $C=C(\eta,\alpha)$ such that if the system is $\eta$-spectral independence then the spectral gap of the $\alpha n$-uniform block dynamics is $\gamma(\mathrm{P}_{n,(1-\alpha)n}^{\vee})\geq C$.
@@
$\Gamma_{(1-\alpha/2)n}=\prod_{i=0}^{(1-\alpha/2)n-2}(2\gamma_{i}-1)\\ \geq\prod_{i=0}^{(1-\alpha/2)n-1}\left(1-\frac{2\eta}{n-i-1}\right)\geq\left(1-\frac{4\eta}{\alpha n}\right)^{n}\geq\exp(-8\eta/\alpha),$ (53)
@@ 6.4 Shattering
The first step is to look at the properties of the updated vertices in a step of the block dynamics. In particular, [CLV21a] showed that a random subset of $\alpha n$ vertices is “shattered” in the sense that the expected size of each component is $O(1)$ when $\alpha<1/(6\Delta)$ where $\Delta$ is the maximum degree of the graph $G$. The shattering occurs because $\alpha\Delta<1$ and hence this corresponds to a sub-critical branching process.
@@
[CLV21a, Lemma 4.3] For a graph $G=(V,E)$ of maximum degree $\Delta$, for $\alpha>0$, choose a random subset $S\subset V$ where $|S|=\alpha n$. Then for every integer $k\geq 1$, for every $v\in V$,
$\mathrm{Pr}\left(|T_{v}|=k\right)\leq\alpha(6\Delta\alpha)^{k-1},$
where $T_{v}$ is the component containing $v$ in the induced subgraph on $S$.
@@
The number of distinct subgraphs of size $k$ that contain $v$ is bounded by $\binom{\Delta k}{k-1}$ (we can enccode them by a binary string of length $\Delta k$ with exactly k1k-1 ones as follows: running the DFS algorithm starting from $v$, for each encountered edge output $1$ if it is traversed and 0 otherwise, and at the end pad the string with zeros). The probability that a particular set of size $k$ is contained in $S$ is $\binom{n-k}{\alpha n-k}/\binom{n}{\alpha n}\leq\alpha^{k}$. Combining the two estimates with $\binom{n}{k}\leq(n{\mathrm{e}}/k)^{k}$ we obtain
$\mathrm{Pr}\left(|T_{v}|=k\right)\leq\left(\frac{\Delta{\mathrm{e}}k}{k-1}\right)^{k}\alpha^{k}\leq\alpha(6\Delta\alpha)^{k-1}.$
6.5 Optimal Relaxation Time of Glauber: Proof of Theorem 1.5
Using the above shattering result (Lemma 6.3) with the fast mixing result for the $\alpha n$-uniform block dynamics (Lemma 6.2), we can prove optimal upper bound on the relaxation time of the Glauber dynamics and hence prove Theorem 1.5.
The high-level proof idea for Theorem 1.8 is the following. Let $S\subset V$ be a random subset where $|S|=\alpha n$ as in Lemma 6.3. Let $\mathcal{C}_{S}$ denote the connected components in the induced subgraph on $S$. When applying the heat-bath dynamics on $S$, then each $T,T^{\prime}\in\mathcal{C}_{S}$ are independent of each other. Hence, given a configuration $\tau$ on $\overline{S}$, then since we have a product distribution over each component $T\in\mathcal{C}_{S}$ we have:
$\mathrm{Var}_{S}[F|\tau]=\sum_{T\in\mathcal{C}_{S}}\mathrm{Var}_{T}[F|\tau].$
Within a component $T\in\mathcal{C}_{S}$ we can apply Theorem 1.4 to conclude a polynomial bound on the spectral gap of the Glauber dynamics within $T$ (with fixed configuration $\tau$ on the boundary of $T$), and hence
$\mathrm{Var}_{T}[F|\tau]\leq C^{\prime}|T|^{\eta+1}\sum_{v\in T}\mathcal{E}_{\mathrm{P^{\tau}_{Gl}}(T)}(f).$ (54)
Finally, we apply Lemma 6.3 to utilize the upper bound on the size of $T$ and conclude the proof.
The bound in Eq. 54 is the only step in the proof where the analogous bound does not hold with respect to entropy in place of variance. This is an issue when we aim to upper bound the mixing time instead of the relaxation time. To obtain a bound on the entropy analog of Eq. 54 we need the additional assumption of marginal boundedness; see Section 6.8 for further details.
Proof of Theorem 1.5.
@@
$\mathcal{E}_{\mathrm{P_{\textsc{gd}}}}(f)$ $=\frac{1}{2}\sum_{\tau\in\Omega}\sum_{\sigma\in\Omega}\mu(\tau)\mathrm{P_{\textsc{gd}}}(\tau,\sigma)(f(\tau)-f(\sigma))^{2}$
$=\sum_{\tau\in\Omega}\mu(\tau)\frac{1}{n}\sum_{v\in V}\left[\frac{1}{2}\sum_{\begin{subarray}{c}\sigma\in\Omega:\\ \sigma(V\setminus\{v\})=\tau(V\setminus\{v\})\end{subarray}}\mu(\sigma)(f(\tau)-f(\sigma))^{2}\right]$
@@
$=\frac{1}{C}\frac{1}{\binom{n}{\alpha n}}\sum_{\begin{subarray}{c}S\subset V:\\ |S|=\alpha n\end{subarray}}\sum_{\tau\in\Omega}\mu(\tau)\mathcal{E}_{\mathrm{P_{HB}^{\tau}}(S)}(f)$ (see below)
$=\frac{1}{C}\frac{1}{\binom{n}{\alpha n}}\sum_{\begin{subarray}{c}S\subset V:\\ |S|=\alpha n\end{subarray}}\sum_{\tau\in\Omega}\mu(\tau)\mathrm{Var}_{S}[F\mid\tau]$ by Eq. 55
$=\frac{1}{C}\frac{1}{\binom{n}{\alpha n}}\sum_{\begin{subarray}{c}S\subset V:\\ |S|=\alpha n\end{subarray}}\sum_{\tau\in\Omega}\mu(\tau)\sum_{T\in\mathcal{C}_{S}}\mathrm{Var}_{T}[F\mid\tau]$ by independence of $T,T^{\prime}\in\mathcal{C}_{S}$
$\leq\frac{1}{C}\frac{1}{\binom{n}{\alpha n}}\sum_{\begin{subarray}{c}S\subset V:\\ |S|=\alpha n\end{subarray}}\sum_{\tau\in\Omega}\mu(\tau)\sum_{T\in\mathcal{C}_{S}}C^{\prime}|T|^{\eta+1}\sum_{v\in T}\mathcal{E}_{\mathrm{P^{\tau}_{Gl}}(T)}(f)$ by Theorem 1.4, (see below)
$=\frac{1}{C}\frac{1}{\binom{n}{\alpha n}}\sum_{\begin{subarray}{c}S\subset V:\\ |S|=\alpha n\end{subarray}}\sum_{\tau\in\Omega}\mu(\tau)\sum_{T\in\mathcal{C}_{S}}C^{\prime}|T|^{\eta+1}\frac{1}{|T|}\sum_{v\in T}\mathrm{Var}_{v}[F\mid\tau]$ by Eq. 56
$=\frac{1}{C}\sum_{\tau\in\Omega}\mu(\tau)\sum_{v\in V}\mathrm{Var}_{v}[F\mid\tau]\sum_{k=1}^{\alpha n}\mathrm{Pr}\left(|T_{v}|=k\right)\times C^{\prime}k^{\eta-1}$ rearranging
@@
$=\frac{n}{C^{\prime\prime}}\mathcal{E}_{\mathrm{P_{\textsc{gd}}}}(f)$ by Eq. 56.
In the second line we are implementing the $(n,(1-\alpha)n)$ down-up dynamics by first choosing the subset $S$ of size $\alpha n$ for update and then applying the heat-bath dynamics on this set $S$ with those vertices outside $S$ fixed to $\tau$ chosen from the Gibbs distribution $\mu$. And in the fifth line where we are applying Theorem 1.4, note we are considering the Glauber dynamics which only operates on the vertices in this component $T$, where the vertices outside $T$ have the fixed configuration $\tau$.
This proves that the spectral gap of the Glauber dynamics is $\geq C^{\prime\prime}/n$ for a constant $C^{\prime\prime}=C^{\prime\prime}(\alpha,\Delta)$. Hence, by applying Eq. 16, the relaxation time of the Glauber dynamics is $\leq Cn$ for some constant $C=C(\alpha,\Delta)$. For the mixing time bound, since $\log(1/\mu^{*})=O(n\log{q})$ and we are considering the case $q=2$, then by Eq. 16, we have $T_{\mathrm{mix}}\leq Cn^{2}$ for some constant $C=C(\alpha,\Delta)$. This establishes the bounds on the relaxation time and mixing time stated in Theorem 1.5. ∎
6.6 Improved Random Walk Theorem: Proof of Theorem 6.1
In this section we will prove the improved Random Walk Theorem presented in Theorem 6.1 in Section 6.2. The proof will involve statements comparing Dirichlet forms and variance on different levels and hence we will need to define the projection of a function $f$ on lower levels.
@@
We defer the proofs of Lemmas 6.4 and 6.5 to Section 6.7, and we proceed with the proof of Theorem 6.1.
Proof of Theorem 6.1.
We begin by proving Eq. 51 which states that the spectral gap of $\mathrm{P}_{k}^{\vee}$ is at least $\Gamma_{k-1}/(\sum_{i=0}^{k-1}\Gamma_{i}$ where $\Gamma_{i}=\prod_{j=0}^{i-1}(2\gamma_{j}-1)$. We will then use this result for the one-level down-up walk to derive the more general statement in Theorem 6.1 for the down-up walk $\mathrm{P}_{n,\ell}^{\vee}$.
@@
$=\Gamma_{k}\mathcal{E}_{\mathrm{P}_{k+1}^{\vee}}(f^{(k+1)})+\Gamma_{k}\mathrm{Var}_{\pi_{k}}(f^{(k)})$ $\mbox{since }\Gamma_{k}=\gamma_{k-1}\Gamma_{k-1}$
@@
Finally, applying Lemma 6.4 with $i=n,j=\ell$, we have that Eq. 62 is equivalent to Eq. 60, which completes the proof of Theorem 6.1. ∎